# Finite Elements with weighted bases for the fractional Laplacian

### v1

@@ 2.1 Proof of convergence rates in $H^{s}$
$[v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}$ $\leq\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}\left[E_{\alpha}r_{\Omega}[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right]_{H^{s}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\alpha}r_{\Omega}[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right\|_{L^{\infty}(\mathbb{R}^{d})}$
$\lesssim\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\alpha}r_{\Omega}[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right\|_{W^{\alpha,\infty}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)}$
$\lesssim\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)}$
$\leq\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)},$
@@
$\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)}\lesssim h^{\mu-\alpha}\left\|E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)}\simeq h^{\mu-\alpha}\|\frac{v}{\delta^{s}}\|_{W^{\mu,\infty}({\Omega})}.$
@@
We are now ready to proof our main result of convergence.
@@ 3 Extensions and open problems
Throughout the paper we work with domains $\Omega$ that are, at least, $C^{1}$. This regularity is used to apply the Hölder regularity theory of $u/\delta^{s}$, to prove $\delta^{s}\in H^{s}_{0}(\Omega)$ (in particular to show $\delta^{s-1}\in L^{1}(\Omega)$), and for the existence suitable extension operator in $W^{1,\infty}(\Omega)$. The latter technical facts probably hold under less regularity of $\Omega$, and hence Theorem 1.7 can be sharpened. However, to weaken the regularity assumption of $\Omega$ in Corollary 1.8 would require improving the regularity theory of $u/\delta^{s}$.
@@ 4 Numerical Experiments
We know present numerical experiments in one dimension. In order to compute the entries of the stiffness matrix $A_{i,j}=a(\phi_{j},\phi_{i})$ we have used a quadrature of the singular integral, that we will discuss in an upcoming work [22]. The experiments have been implemented in julia 1.11.7, and the codes are available at
@@ 4.2 $H^{s}$ and $L^{2}$ rates when $f=1$
As before, let $u^{*}$ be the explicit solution in (1.11) with $d=1$, $R=1$, and $x_{0}=0$, i.e., $\Omega=(-1,1)$ and $f=1$. In Figure 4 we validate the theoretical results that we have presented in Theorems 1.7, 1.8 and 1.9. In particular, with the choice of $f=1$ we put ourselves in the case of maximal $H^{s}$-rate of Corollary 1.8, i.e., $h^{2-s}\lvert\log h\rvert^{\frac{1}{2}}$. As shown in the Figure 44(a) the observed convergence rates match the theoretical prediction, with the observed slope of every convergence line in log scale for different $s$ being approximately 2s2-s. Figure 44(b) instead is linked to the fact that, even we are only able to prove Theorem 1.9, we would expect the convergence to be quadratic in $L^{2}(\Omega)$.

### v2

@@ 2.1 Proof of convergence rates in $H^{s}$
$[v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}$ $\leq\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}\left[E_{\alpha}r_{\Omega}\normalcolor[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right]_{H^{s}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\alpha}r_{\Omega}\normalcolor[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right\|_{L^{\infty}(\mathbb{R}^{d})}$
$\lesssim\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\alpha}r_{\Omega}\normalcolor[E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}]\right\|_{W^{\alpha,\infty}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)\normalcolor}$
$\lesssim\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega\normalcolor)}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)\normalcolor}$
$\leq\frac{\|\delta^{s}\|_{L^{\infty}(\mathbb{R}^{d})}+[\delta^{s}]_{H^{s}(\mathbb{R}^{d})}}{\sqrt{\alpha-s}}\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega\normalcolor)},$
@@
$\left\|E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)\normalcolor}\lesssim h^{\mu-\alpha}\left\|E_{\mu}\frac{v}{\delta^{s}}\right\|_{W^{\alpha,\infty}(\Omega)\normalcolor}\simeq h^{\mu-\alpha}\|\frac{v}{\delta^{s}}\|_{W^{\mu,\infty}({\Omega})}.$
@@
We are now ready to prove our main result of convergence.
@@ 3 Extensions and open problems
Throughout the paper we work with domains $\Omega$ that are, at least, $C^{1,1}$. This regularity is used to apply the Hölder regularity theory of $u/\delta^{s}$, to prove $\delta^{s}\in V$ (in particular to show $\delta^{s-1}\in L^{1}(\Omega)$), and for the existence suitable extension operator in $W^{1,\infty}(\Omega)$. The latter technical facts probably hold under less regularity of $\Omega$, and hence Theorem 1.9 can be sharpened. However, to weaken the regularity assumption of $\Omega$ in Corollary 1.10 would require improving the regularity theory of $u/\delta^{s}$.
@@ 4 Numerical Experiments
We now present numerical experiments in one dimension. In order to compute the entries of the stiffness matrix $A_{i,j}=a(\phi_{j},\phi_{i})$ we have used a quadrature of the singular integral, that we will discuss in an upcoming work [25]. The experiments have been implemented in julia 1.11.7, and the codes are available at
@@ 4.2 $H^{s}$ and $L^{2}$ rates when $f=1$
As before, let $u^{*}$ be the explicit solution in (1.10) with $d=1$, $R=1$, and $x_{0}=0$, i.e., $\Omega=(-1,1)$ and $f=1$. In Figure 4 we validate the theoretical results that we have presented in Theorems 1.9, 1.10 and 1.11. In particular, with the choice of $f=1$ we put ourselves in the case of maximal $H^{s}$-rate of Corollary 1.10, i.e., $h^{2-s}\lvert\log h\rvert^{\frac{1}{2}}$. As shown in the Figure 44(a) the observed convergence rates match the theoretical prediction, with the observed slope of every convergence line in log scale for different $s$ being approximately 2s2-s. Figure 44(b) instead is linked to the fact that, even we are only able to prove Theorem 1.11, we would expect the convergence to be quadratic in $L^{2}(\Omega)$.
@@
In Table 1 we report the observed convergence rates for the example corresponding to the slope of the segments in Figure 44(a). The values have been computed using the incremental ratio formula $(\log[e_{h}]_{H^{s}({\mathbb{R}^{d}})}-[e_{2h}]_{H^{s}({\mathbb{R}^{d}})})/(\log h-\log 2h)$, where we recall that $e_{h}=u_{h}-u^{\ast}$.